OpenStudy (anonymous):

Please help i will give a medal and fan : Find the prime factors of the following numbers. Explain each number as a product of prime number using indices when needed. a) 2310 b) 1440

OpenStudy (danjs):

since they are even, start by dividing by 2

OpenStudy (danjs):

2310 = 2 x 1155

OpenStudy (anonymous):

ok

OpenStudy (danjs):

A number ending in 5 will be divisible by 5 2310 = 2 * 5 * 231

OpenStudy (anonymous):

ok

OpenStudy (danjs):

231 is divisible by 3, take out a 3 2310 = 2 x 5 x 3 x 77

OpenStudy (danjs):

77 is divisible by 11 \[\large\color{blue}{ 2310 = 2 *5*3*11*7}\]

OpenStudy (danjs):

that is the prime factorization of 2310

OpenStudy (danjs):

2310 = 2 x 3 x 5 x 7 x 11

OpenStudy (danjs):

FOR 1440, what would you do first?

OpenStudy (danjs):

First thing i do , if it is even, keep pulling out 2's as many times as possible

OpenStudy (danjs):

1440 = 720 x 2

OpenStudy (danjs):

1440 = 360 x 2 x 2

OpenStudy (danjs):

1440 = 180 x 2 x 2 x 2

OpenStudy (danjs):

1440 = 90 x 2 x 2 x 2 x 2

OpenStudy (danjs):

45 x 2 x 2 x 2 x 2 x 2

OpenStudy (danjs):

9 x 5 x 2^5

OpenStudy (danjs):

3 x 3 x 5 x 2^5

OpenStudy (danjs):

\[\large\color{blue}{ 1440 = 2^5*3^2 *5}\]

OpenStudy (triciaal):